OpenStudy (anonymous):

A particle moves along a straight line with velocity v(t) = te^-t meters per second after t seconds. How far will it travel during the first 75 seconds?

sam (.sam.):

\[ v(t) = te^{-t} \] \[\frac{dx}{dt}=te^{-t}\] \[\int\limits_{}^{}1dx=\int\limits_{}^{} te^{-t}dt\] \[x=\int\limits\limits_{}^{} te^{-t}dt\]

OpenStudy (anonymous):

umm

sam (.sam.):

integration by parts

OpenStudy (anonymous):

do the x= part using IBP?

sam (.sam.):

\[x=\int\limits_{0}^{75} e^{-t} t \, dt\] \[x\text{}=\int\limits\limits_{0}^{75} e^{-t} \, dt-e^{-t} t\] \[\large x=[-e^{-t} t-e^{-t}]_0^{75} \]

OpenStudy (anonymous):

ok, just going by your last post.. i don't get the second step.

OpenStudy (anonymous):

i dont get the -et^-t part

OpenStudy (anonymous):

isn't that the integral?... not sure why you stuck it on as a subtraction

sam (.sam.):

Its from integration by parts, \[x=\int\limits\limits_{0}^{75} e^{-t} t \, dt\] let, u=t dv=e^-t du=dt v=-e^-t \[x=-te^{-t}-\int\limits_{}^{}-e^{-t}dt\] Rearranging it \[x\text{}=\int\limits\limits\limits_{0}^{75} e^{-t} \, dt-e^{-t} t\]

OpenStudy (anonymous):

ok, it's starting to make sense

OpenStudy (anonymous):

let me look it over for a bit

OpenStudy (anonymous):